# MCP Bridge: A Lightweight, LLM-Agnostic RESTful Proxy for Model Context Protocol Servers

### v2

@@ Abstract
Large Language Models (LLMs) are increasingly augmented with external tools through standardized interfaces like the Model Context Protocol (MCP). However, current MCP implementations face critical limitations: they typically require local process execution through STDIO transports, making them impractical for resource-constrained environments like mobile devices, web browsers, and edge computing. We present MCP Bridge, a lightweight RESTful proxy that connects to multiple MCP servers and exposes their capabilities through a unified API. Unlike existing solutions, MCP Bridge is fully LLM-agnostic, supporting any backend regardless of vendor. The system implements a risk-based execution model with three security levels—standard execution, confirmation workflow, and Docker isolation—while maintaining backward compatibility with standard MCP clients. However, reliable execution within this framework requires models that can strictly adhere to protocol schemas. To this end, we also fine-tuned the Qwen3 4B and 8B model family on the Agent-Ark/Toucan-1.5M dataset using four Reinforcement Learning techniques: Group Relative Policy Optimization (GRPO), Dr. GRPO, Beta Normalization Policy Optimization (BNPO), and Decoupled Alignment Policy Optimization (DAPO). Evaluated on the MCPToolBench++ benchmark, our optimized model achieves an F1 score of 73.0% that outperforms GPT-OSS-120B (62.17%) and remains competitive with the 70B+ parameter baselines. Evaluation demonstrates that MCP Bridge successfully addresses the constraints of direct MCP connections while providing enhanced security controls and cross-platform compatibility, enabling sophisticated LLM-powered applications in previously inaccessible environments.
@@ 3.5 Client Integration (MCP-Gemini Agent)
Beyond demonstrating MCP Bridge with a proprietary LLM client, the proxy is designed to be model-agnostic: any language model capable of emitting MCP-style structured tool calls can use the same REST endpoints to discover tools, invoke them, and receive results. In the next section, we describe how we fine-tune open-weight Qwen3 models to produce MCP-compliant tool calls reliably, enabling them to act as drop-in clients of MCP Bridge.
@@ 4.6 Implementation Details (Training Configuration)
We implement training using Unsloth [19] for efficient Qwen3 loading and LoRA adaptation, and TRL-style policy optimization with group sampling. Table 2 summarizes the configuration used in our training script.
@@ 5 Experiments and Results
We evaluate tool-use behavior on MCPToolBench++ [10], reporting Precision, Recall, F1, and Accuracy for six tool categories (Browser, File System, Search, Map, Pay, Finance). We compare the base Qwen3 models against four RL methods (GRPO, Dr. GRPO, DAPO, BNPO), trained on tool-use data derived from Toucan-1.5M [5] using the tool-selection and format rewards described in Section 4.
@@ 5.1 Per-category F1 trends
A radar view of the same F1 values is shown in Figure 4. For Qwen3-4B, Dr. GRPO improves overall balance across Browser, Search, and Pay relative to the base model, whereas DAPO shows mixed gains with a notable drop in Finance. For Qwen3-8B, BNPO expands the performance most strongly in Browser while maintaining competitive Pay and Search, while Dr. GRPO produces the highest Finance score.
@@ 5.2 Aggregate metrics and method comparison
Figure 5 reports aggregate Precision, Recall, F1, and Accuracy (averaged over categories). For Qwen3-4B, Dr. GRPO yields the strongest aggregate F1 and Accuracy, improving tool selection reliability relative to the base model while preserving high File System performance. GRPO is close behind, indicating that group-relative updates alone deliver meaningful gains when the reward signal is dominated by selection correctness and format validity. In contrast, DAPO and BNPO do not consistently outperform Dr. GRPO at 4B, suggesting that small models are more sensitive to reward distribution shifts and clipping choices, particularly in Finance.
For Qwen3-8B, BNPO achieves the highest aggregate Precision/Recall/F1 and the best Accuracy, indicating that as capacity increases, bounded/normalized updates can convert the same reward signal into more consistent improvements across categories. Notably, Finance improves substantially for 8B under GRPO-family methods, implying that higher-capacity policies can better internalize tool-call patterns and disambiguate tool choice even when category semantics are difficult. An ablation study isolating the tool-selection reward and the format reward is provided in Appendix A.
Figure 5: Aggregate Precision, Recall, F1, and Accuracy (%) on MCPToolBench++ [10] for Qwen3-4B (top) and Qwen3-8B (bottom). Scores are averaged across the six tool categories. These results show that Dr. GRPO is the strongest overall method for Qwen3-4B, while BNPO is strongest for Qwen3-8B under the same reward design.
@@ 5.3 Comparison against other models
Table 3 compares our best MCP-aligned checkpoints to several representative baselines using the same metric suite. The comparison contextualizes the tool-use reliability of MCP-aligned Qwen3 models against larger models that may have broader general capabilities. While the largest baselines achieve higher absolute scores, the RL-aligned Qwen3 models narrow the gap substantially relative to their base versions, demonstrating that targeted policy optimization with selection+format rewards can produce practical tool-use behavior at smaller scales.
Table 3: Tool-use evaluation comparison (Precision/Recall/F1/Accuracy).
Model Precision (%) Recall (%) F1 (%) Accuracy (%)
MCP Bridge (Qwen3-4B + Dr. GRPO) 66.0 66.7 66.2 65.3
MCP Bridge (Qwen3-8B + BNPO) 72.2 74.7 73.0 69.7
GPT-OSS-120B 61.11 64.33 62.17 58.00
Kimi-K2-Instruct 76.83 80.00 77.89 73.67
Llama-3.3-70B 81.50 82.00 81.67 81.00
Llama-4-Maverick-17BA-400B 78.18 80.67 78.94 76.00
Llama-4-Scout-17BA-109B 76.36 77.67 76.67 75.67
@@ 7 Conclusion
Beyond the proxy layer, we showed that reliable MCP usage depends on the client model’s ability to produce tool calls that are both correct and protocol-compliant. To support open-weight deployment, we fine-tuned Qwen3-4B and Qwen3-8B using four RL methods (GRPO, Dr. GRPO, DAPO, BNPO) on Toucan-1.5M and evaluated them on MCPToolBench++. Across model sizes, RL improves tool-use reliability relative to the base models, with Qwen3-4B + Dr. GRPO achieving 66.2% F1 and 65.3% Accuracy, and Qwen3-8B + BNPO achieving 73.0% F1 and 69.7% Accuracy. These results outperform the GPT-OSS-120B baseline in our comparison table and narrow the gap to much larger 70B+ parameter baselines.
@@ Appendix A Reward Ablation Study
This appendix isolates the effect of the two reward components used in our RL training: the tool-selection reward and the format reward. We compare two ablations—Tool Select. Only and Format Only—against the main setting used in Section 5 (tool-selection + format). Across both Qwen3-4B and Qwen3-8B, the ablations show that the two reward components are complementary: optimizing formatting alone can improve protocol compliance and yield moderate metric gains, while optimizing selection alone provides limited benefit and can be unstable for smaller models. The best overall performance is achieved when both rewards are used jointly in the main experiments.
Table 4: Summary of best aggregate results (%) under reward ablations (macro-averaged over categories).
Model Full (Sel.+Fmt.) Tool Select. Only Format Only
Qwen3-4B (best F1 / Acc.) 66.2 / 65.3 63.1 / 61.3 65.8 / 64.0
Qwen3-8B (best F1 / Acc.) 73.0 / 69.7 71.6 / 68.7 72.3 / 68.3

### v3

@@ Abstract
Large Language Models (LLMs) are increasingly augmented with external tools through standardized interfaces like the Model Context Protocol (MCP). However, current MCP implementations face critical limitations: they typically require local process execution through STDIO transports, making them impractical for resource-constrained environments like mobile devices, web browsers, and edge computing. We present MCP Bridge, a lightweight RESTful proxy that connects to multiple MCP servers and exposes their capabilities through a unified API. Unlike existing solutions, MCP Bridge is fully LLM-agnostic, supporting any backend regardless of vendor. The system implements a risk-based execution model with three security levels—standard execution, confirmation workflow, and Docker isolation—while maintaining backward compatibility with standard MCP clients. However, reliable execution within this framework requires models that can strictly adhere to protocol schemas. To this end, we also fine-tuned the Qwen3 4B and 8B model family on the Agent-Ark/Toucan-1.5M dataset using four Reinforcement Learning techniques: Group Relative Policy Optimization (GRPO), Dr. GRPO, Beta Normalization Policy Optimization (BNPO), and Decoupled Clip and Dynamic sAmpling Policy Optimization (DAPO). Evaluated on the MCPToolBench++ benchmark, our optimized model achieves an F1 score of 73.0% that outperforms GPT-OSS-120B (62.17%) and remains competitive with the 70B+ parameter baselines. Evaluation demonstrates that MCP Bridge successfully addresses the constraints of direct MCP connections while providing enhanced security controls and cross-platform compatibility, enabling sophisticated LLM-powered applications in previously inaccessible environments.
@@ 3.5 Client Integration (MCP-Gemini Agent)
Beyond demonstrating MCP Bridge with a proprietary LLM client, the proxy is designed to be model-agnostic: any language model capable of emitting MCP-style structured tool calls can use the same REST endpoints to discover tools, invoke them, and receive results. In the next section, we present an empirical evaluation of MCP Bridge’s system-level performance before describing how we fine-tune open-weight Qwen3 models for MCP tool alignment.
3.6 System Performance Evaluation
To validate the architectural claims presented above and quantify the performance characteristics of the REST proxy layer, we conducted a systematic benchmark of MCP Bridge measuring end-to-end latency, throughput under concurrent load, risk-level execution-path overhead, and resource utilization. All measurements were collected on a single machine running Windows 11 with an AMD Ryzen 7 7435HS (16 logical cores) and 24,261 MB RAM, using Node.js v22.12.0. To ensure statistical reliability, latency and concurrency tests were repeated over three independent runs with 50 iterations per operation per run; we report mean $\pm$ standard deviation across runs.
3.6.1 Benchmark Configuration
Four MCP servers were started simultaneously through MCP Bridge using the configuration summarized in Table 2. The servers span three distinct MCP server implementations and two risk levels, exercising 50 tools in total.
Table 2: MCP Bridge benchmark configuration. Four STDIO-based MCP servers were started simultaneously, spanning three distinct server implementations and two risk levels.
Server ID Implementation Risk Level Tools
filesystem @modelcontextprotocol/server-filesystem 1 (Low) 14
filesystem-medium @modelcontextprotocol/server-filesystem 2 (Medium) 14
memory @modelcontextprotocol/server-memory 1 (Low) 9
everything @modelcontextprotocol/server-everything 1 (Low) 13
Total tools discovered 50
3.6.2 Latency Analysis
We measured round-trip latency for six representative tool-execution operations under three access modes: (i) Bridge REST, where the client sends an HTTP request to MCP Bridge and the proxy forwards it to the MCP server over STDIO; (ii) STDIO keep-alive, a direct STDIO connection that remains open across requests (best-case baseline); and (iii) STDIO per-spawn, where a new server process is spawned for every request (worst-case baseline). Table 3 reports the results.
Table 3: Tool-execution latency comparison (ms). Bridge REST values are mean $\pm$ std across 3 runs of 50 iterations each. STDIO values are from a single run (50 iterations keep-alive, 10 iterations per-spawn). Bridge overhead is Bridge REST mean minus STDIO keep-alive mean.
Operation Bridge REST STDIO Keep-Alive STDIO Per-Spawn Overhead
filesystem/list_directory 2.55 $\pm$ 0.35 0.91 5.60 1.64
filesystem/read_file 2.44 $\pm$ 0.03 1.10 6.72 1.34
everything/echo 1.55 $\pm$ 0.10 0.48 5.24 1.07
everything/get-sum 1.66 $\pm$ 0.08
memory/create_entities 2.98 $\pm$ 0.08
memory/read_graph 2.15 $\pm$ 0.08 0.80 5.03 1.35
[…]
3.6.3 Throughput Under Concurrent Load
To evaluate scalability, we measured throughput (requests/second) and mean latency at concurrency levels of 1, 5, 10, 20, and 50 simultaneous clients. Each level issued 20–100 mixed tool-execution requests distributed across the four servers. Table 4 and Figure 4 report the results.
Table 4: Throughput and latency under concurrent load (mean $\pm$ std, 3 runs). MCP Bridge maintains zero errors across all concurrency levels.
Concurrency Throughput (req/s) Mean Latency (ms) Total Requests Errors
1 430 $\pm$ 31 2.26 $\pm$ 0.15 20 0
5 827 $\pm$ 19 5.08 $\pm$ 0.10 100 0
10 839 $\pm$ 46 10.47 $\pm$ 0.53 100 0
20 920 $\pm$ 61 18.20 $\pm$ 1.05 100 0
50 909 $\pm$ 88 43.14 $\pm$ 3.63 100 0
Throughput scales from 430 req/s (single client) to over 900 req/s at 50 concurrent clients, demonstrating that the asynchronous Node.js architecture effectively exploits I/O concurrency. Latency increases approximately linearly with concurrency, which is expected for a single-threaded event-loop proxy, while the error rate remains at 0% across all levels.
Figure 4: Throughput scaling as a function of concurrent clients (mean $\pm$ std, 3 independent runs). MCP Bridge doubles throughput from 1 to 5 clients and maintains >>900 req/s at 50 concurrent clients with zero errors.
3.6.4 Risk-Level Execution Traces
To validate each execution path described in Section 3.4, we recorded end-to-end latency traces for the three risk levels using real MCP server requests.
  • Level 1 (Direct): A list_directory call to the filesystem server (risk level 1) completed in 4.61 ms with a single HTTP round trip.
  • Level 2 (Confirmation): The same tool on filesystem-medium (risk level 2) required two round trips: the initial request returned a confirmation token in 1.35 ms, and execution after confirmation completed in 6.82 ms, totaling 8.16 ms. Re-submitting the same token returned HTTP 404, confirming single-use enforcement.
  • Level 2 (Rejection): Submitting a rejection for a pending confirmation correctly returned a cancellation response without executing the tool.
Figure 5 shows the latency breakdown. The confirmation workflow adds approximately 3.6 ms of overhead compared to direct execution, a modest cost for the additional safety guarantee.
Figure 5: Risk-level execution-path latency. Level 1 (direct) completes in a single round trip; Level 2 (confirmation) requires two round trips with a single-use token validator.
3.6.5 Resource Utilization and Cold Start
Figure 6 shows process memory and system free memory during a sustained-load test (200 requests at concurrency 10). The benchmark process RSS remains stable at $\approx$47 MB and heap usage at $\approx$7–8 MB, indicating no memory leaks. System free memory ($\approx$8.9 GB) is unaffected, confirming a minimal resource footprint.
Cold-start time—measured by launching MCP Bridge with all four servers from a stopped state—was 8,722 ms, dominated by Node.js module loading and MCP server initialization. Once started, subsequent tool calls incur only the REST latency described above. Additional benchmark figures (per-operation distributions, percentile heatmaps, CDF curves, API endpoint latency, and cold-start breakdown) are provided in Appendix C.
Figure 6: Resource utilization during a sustained-load test (200 requests, concurrency 10). Process RSS and heap usage remain stable, and system free memory is unaffected, confirming a minimal resource footprint under load.
3.6.6 Bridge Overhead Analysis
Figure 7 isolates the REST proxy overhead by subtracting STDIO keep-alive latency from Bridge REST latency for each operation. The mean overhead across operations is 1.35 ms (range: 1.07–1.64 ms), confirming that the HTTP/JSON serialization layer adds negligible cost relative to the convenience of network-accessible, multi-client MCP access.
Figure 7: Per-operation REST proxy overhead (Bridge REST minus STDIO keep-alive). The average overhead is 1.35 ms, attributable to HTTP parsing, JSON serialization, and Express.js routing.
@@ 4.6 Implementation Details (Training Configuration)
We implement training using Unsloth [19] for efficient Qwen3 loading and LoRA adaptation, and TRL-style policy optimization with group sampling. All training runs and ablation experiments were carried out on a single NVIDIA RTX 6000 Pro GPU (96 GB VRAM). Table 5 summarizes the configuration used in our training script.
@@ 5 Experiments and Results
We evaluate tool-use behavior on MCPToolBench++ [10], reporting Precision, Recall, F1, and Accuracy for six tool categories (Browser, File System, Search, Map, Pay, Finance), with 50 randomly sampled evaluation instances per category (N=300N{=}300 total; see Table 8). We compare the base Qwen3 models against four RL methods (GRPO, Dr. GRPO, DAPO, BNPO), trained on tool-use data derived from Toucan-1.5M [5] using the tool-selection and format rewards described in Section 4. All reported 95% confidence intervals (CIs) are computed via 10,000-iteration bootstrap resampling of per-sample scores. [^1] All per-sample evaluation results were regenerated from scratch to compute bootstrap confidence intervals; the underlying models and test set are unchanged.
@@ 5.1 Per-category F1 trends
A radar view of the same F1 values is shown in Figure 9. For Qwen3-4B, GRPO yields the largest gains, improving Browser from 70% to 84% and Pay from 72.7% to 83.3%; Dr. GRPO follows with the highest Search gain (66% vs. 62% base) and the best Pay (86.7%). DAPO and BNPO show more moderate overall gains. For Qwen3-8B, Dr. GRPO produces the strongest result, raising Search from 54% to 68% and maintaining Finance (84%); DAPO is close behind with the highest Browser (70.7%), while GRPO achieves the best Finance (90%) at a cost to Map and Pay.
@@ 5.2 Aggregate metrics and method comparison
Figure 10 reports aggregate Precision, Recall, F1, and Accuracy (averaged over categories); full per-variant results with 95% CIs are in Table 10. For Qwen3-4B, GRPO yields the strongest aggregate F1 (67.4%, 95% CI [62.2, 72.7]) among models trained with both rewards, improving tool selection reliability relative to the base model (F1 62.7%, CI [57.2, 68.0]). Dr. GRPO is close behind (F1 66.7%), indicating that reference-distribution regularization also delivers meaningful gains when the reward signal is dominated by selection correctness and format validity.
For Qwen3-8B, Dr. GRPO achieves the highest aggregate F1 (73.4%, CI [68.6, 78.2]) and the best Accuracy (69.7%), indicating that as capacity increases, reference-aware updates can convert the same reward signal into more consistent improvements across categories.
A failure-type analysis of the base models (Table 7) sheds light on what RL training addresses. For the 4B base model, the dominant failure mode is “no tool call” (21.7% of samples), where the model does not produce any parseable tool invocation. At 8B, this drops to 12.3%, but “wrong tool” errors increase to 12.7% (vs. 9.0% at 4B), suggesting that larger models are better at producing well-formed tool calls but still struggle with tool disambiguation. An ablation study isolating the tool-selection reward and the format reward is provided in Appendix A.
Figure 10: Aggregate Precision, Recall, F1, and Accuracy (%) on MCPToolBench++ [10] for Qwen3-4B (top) and Qwen3-8B (bottom). Scores are averaged across the six tool categories. GRPO is the strongest overall method for Qwen3-4B, while Dr. GRPO is strongest for Qwen3-8B under the same reward design.
@@ 5.3 Comparison against other models
Table 6 compares our best MCP-aligned checkpoints to several representative baselines using the same metric suite. All API baselines were served through Groq’s inference API [20], which hosts open-weight models (Llama-3.3-70B, Llama-4-Maverick, Llama-4-Scout, GPT-OSS-120B) as well as the proprietary Kimi-K2-Instruct endpoint. The comparison contextualizes the tool-use reliability of MCP-aligned Qwen3 models against larger models that may have broader general capabilities. While the largest baselines achieve higher absolute scores, the RL-aligned Qwen3 models narrow the gap substantially relative to their base versions, demonstrating that targeted policy optimization with selection+format rewards can produce practical tool-use behavior at smaller scales.
Notably, the 8B Dr. GRPO model achieves higher accuracy (69.7%) than the 120B GPT-OSS-120B baseline (58.7%), with non-overlapping bootstrap CIs, confirming that our 8-billion-parameter RL-aligned model can match or exceed much larger models on this benchmark. The 4B GRPO model (65.7%) likewise outperforms GPT-OSS-120B. While the largest open-weight baselines (e.g. Llama-3.3-70B at 81.0%) still lead, the margins to Kimi-K2, Llama-4-Maverick, and Llama-4-Scout are narrower, indicating that the 8B model is competitive with these mid-range baselines.
Table 6: Tool-use evaluation comparison on MCPToolBench++ (N=300N{=}300; 50 samples $\times$ 6 categories). Values are percentages; 95% CIs from 10,000-iteration bootstrap resampling.
Model Prec. (%) Rec. (%) F1 (%) Acc. (%)
MCP Bridge (4B + GRPO) 67.0 68.3 67.4 [62.2, 72.7] 65.7 [60.3, 71.0]
MCP Bridge (8B + Dr. GRPO) 72.5 75.3 73.4 [68.6, 78.2] 69.7 [64.3, 74.7]
GPT-OSS-120B 62.0 65.3 63.1 [57.8, 68.4] 58.7 [53.0, 64.3]
Kimi-K2-Instruct 77.3 80.7 78.4 [73.9, 82.7] 74.0 [69.0, 78.7]
Llama-3.3-70B 81.8 82.7 82.1 [77.7, 86.2] 81.0 [76.7, 85.3]
Llama-4-Maverick 77.5 80.0 78.3 [73.8, 82.8] 75.3 [70.7, 80.3]
Llama-4-Scout 75.0 76.3 75.3 [70.4, 80.1] 74.3 [69.3, 79.3]
Table 7: Failure-type breakdown for base (un-tuned) Qwen3 models (N=300N{=}300). “No tool call”: no parseable tool invocation; “Wrong tool”: valid call to incorrect tool; “Fmt. heuristic”: output rescued by fallback regex extraction.
Model Correct No tool call Wrong tool Partial Fmt. heuristic
4B Base 62.0% 21.7% 9.0% 1.0% 6.3%
8B Base 68.7% 12.3% 12.7% 4.3% 2.0%
Table 8: Evaluation samples per tool category in our MCPToolBench++ subset.
Category NN
Browser 50
File System 50
Search 50
Map 50
Pay 50
Finance 50
Total 300
@@ 7 Conclusion
Beyond the proxy layer, we showed that reliable MCP usage depends on the client model’s ability to produce tool calls that are both correct and protocol-compliant. To support open-weight deployment, we fine-tuned Qwen3-4B and Qwen3-8B using four RL methods (GRPO, Dr. GRPO, DAPO, BNPO) on Toucan-1.5M and evaluated them on MCPToolBench++ (N=300N{=}300, 50 samples per category). Across model sizes, RL produces consistent directional improvements in tool-use reliability, with Qwen3-4B + GRPO achieving 67.4% F1 (95% CI [62.2, 72.7]) and 65.7% Accuracy, and Qwen3-8B + Dr. GRPO achieving 73.4% F1 (95% CI [68.6, 78.2]) and 69.7% Accuracy. Bootstrap CIs confirm that the 8B model outperforms the GPT-OSS-120B baseline (non-overlapping 95% CIs), although base-to-RL improvements within either model size show overlapping intervals on the current evaluation set.
@@ Appendix A Reward Ablation Study
This appendix isolates the effect of the two reward components used in our RL training: the tool-selection reward and the format reward. We compare two ablations—Tool Select. Only and Format Only—against the main setting used in Section 5 (tool-selection + format). Table 9 summarizes the best aggregate results under each ablation.
Across both Qwen3-4B and Qwen3-8B, the ablations reveal that the two reward components have complementary but asymmetric effects. Optimizing formatting alone can produce sizable gains: for 4B, the format-only ablation (BNPO, F1 67.1%) is close to the full-reward setting (GRPO, F1 67.4%), suggesting that at smaller scales, enforcing output structure alone captures most of the available improvement. For 8B, format-only (GRPO, F1 73.4%) matches the full-reward result (Dr. GRPO, F1 73.4%). In contrast, optimizing tool selection alone yields limited gains for both model sizes (4B: 61.3%, 8B: 68.6%), indicating that selection-only training is insufficient without explicit format incentives. The best overall performance is achieved when both rewards are used jointly, though the margin over format-only is modest.
Table 9: Best aggregate results (%) under reward ablations. Each cell: F1 / Accuracy (best RL method in parentheses).
Model Full (Sel.+Fmt.) Tool Sel. Only Format Only
Qwen3-4B 67.4 / 65.7 (GRPO) 61.3 / 59.3 (DAPO) 67.1 / 65.3 (BNPO)
Qwen3-8B 73.4 / 69.7 (Dr. GRPO) 68.6 / 65.3 (Dr. GRPO) 73.4 / 69.7 (GRPO)
@@
Appendix B Complete Evaluation Results
Table 10 reports the full evaluation results for all 26 local model variants and 5 API baselines on MCPToolBench++ (N=300N{=}300). API baselines were served via Groq [20]. For local models, F1 and Accuracy are reported with 95% bootstrap CIs (10,000 iterations). Rows are grouped by model size and reward configuration, sorted by descending F1 within each group.
Table 10: Complete evaluation results for all model variants on MCPToolBench++ (N=300N{=}300).
Model Prec. (%) Rec. (%) F1 (%) [95% CI] Acc. (%) [95% CI]
Qwen3-4B
     Base 62.5 63.0 62.7 [57.2, 68.0] 62.0 [56.3, 67.3]
     GRPO (Full) 67.0 68.3 67.4 [62.2, 72.7] 65.7 [60.3, 71.0]
     Dr. GRPO (Full) 66.3 67.3 66.7 [61.4, 71.9] 65.3 [60.0, 70.7]
     BNPO (Full) 62.8 64.0 63.2 [57.9, 68.6] 61.7 [56.0, 67.0]
     DAPO (Full) 62.7 63.7 63.0 [57.6, 68.3] 61.7 [56.0, 67.0]
     DAPO (Sel. Only) 60.8 62.3 61.3 [55.8, 66.7] 59.3 [53.7, 64.7]
     GRPO (Sel. Only) 60.2 61.7 60.7 [55.2, 66.0] 58.7 [53.0, 64.3]
     BNPO (Sel. Only) 57.2 58.3 57.6 [52.1, 63.0] 56.0 [50.3, 61.7]
     Dr. GRPO (Sel. Only) 57.2 58.3 57.6 [52.0, 63.0] 56.0 [50.3, 61.7]
     BNPO (Fmt. Only) 66.7 68.0 67.1 [61.9, 72.2] 65.3 [60.0, 70.7]
     Dr. GRPO (Fmt. Only) 64.3 65.7 64.7 [59.4, 70.0] 63.0 [57.7, 68.3]
     DAPO (Fmt. Only) 63.5 65.0 64.0 [58.6, 69.2] 62.0 [56.7, 67.3]
     GRPO (Fmt. Only) 62.7 64.0 63.1 [57.8, 68.3] 61.3 [55.7, 66.7]
Qwen3-8B
     Base 70.8 73.0 71.6 [66.6, 76.4] 68.7 [63.3, 74.0]
     Dr. GRPO (Full) 72.5 75.3 73.4 [68.6, 78.2] 69.7 [64.3, 74.7]
     DAPO (Full) 71.0 73.3 71.8 [66.6, 76.7] 68.7 [63.3, 74.0]
     GRPO (Full) 70.8 73.7 71.7 [66.7, 76.6] 68.0 [62.7, 73.3]
     BNPO (Full) 70.3 72.7 71.1 [66.0, 75.9] 68.0 [62.7, 73.0]
     Dr. GRPO (Sel. Only) 67.8 70.3 68.6 [63.4, 73.8] 65.3 [60.0, 70.7]
     DAPO (Sel. Only) 67.2 69.7 68.0 [62.8, 73.1] 64.7 [59.3, 70.0]
     BNPO (Sel. Only) 67.3 69.0 67.9 [62.7, 73.0] 65.7 [60.3, 71.0]
     GRPO (Sel. Only) 65.6 68.3 66.5 [61.2, 71.7] 63.0 [57.7, 68.3]
     GRPO (Fmt. Only) 72.4 75.3 73.4 [68.4, 78.1] 69.7 [64.3, 74.7]
     DAPO (Fmt. Only) 70.5 73.3 71.4 [66.4, 76.3] 67.7 [62.3, 73.0]
     Dr. GRPO (Fmt. Only) 70.1 72.7 70.9 [65.9, 75.9] 67.7 [62.3, 73.0]
     BNPO (Fmt. Only) 69.8 72.0 70.6 [65.4, 75.6] 67.7 [62.3, 73.0]
API Baselines
     GPT-OSS-120B 62.0 65.3 63.1 [57.8, 68.4] 58.7 [53.0, 64.3]
     Kimi-K2-Instruct 77.3 80.7 78.4 [73.9, 82.7] 74.0 [69.0, 78.7]
     Llama-3.3-70B 81.8 82.7 82.1 [77.7, 86.2] 81.0 [76.7, 85.3]
     Llama-4-Maverick 77.5 80.0 78.3 [73.8, 82.8] 75.3 [70.7, 80.3]
     Llama-4-Scout 75.0 76.3 75.3 [70.4, 80.1] 74.3 [69.3, 79.3]
Appendix C MCP Bridge Benchmark — Supplementary Figures
This appendix provides additional benchmark visualizations that complement the system evaluation in Section 3.6. All measurements were collected under the same configuration described in Table 2 (4 MCP servers, 50 tools, Node.js v22.12.0, AMD Ryzen 7 7435HS, 24 GB RAM).
Figure 17: Bridge REST mean latency per tool-execution operation with error bars (mean $\pm$ std, 3 runs, 50 iterations each). Lightweight operations such as echo and get-sum complete in under 2 ms.
Figure 18: Latency percentile breakdown (P50, P95, P99) per operation. The gap between P50 and P99 is small for most operations, indicating low tail-latency variance.
Figure 19: Latency distributions (min, Q1, median, Q3, max) per operation. The tight interquartile ranges confirm consistent response times across iterations.
Figure 20: Mean latency as a function of concurrent clients (mean $\pm$ std, 3 runs). Latency increases approximately linearly with concurrency, consistent with single-threaded event-loop behavior.
Figure 21: Throughput by concurrency level (bar chart with error bars, 3 runs). Throughput plateaus near 900 req/s at concurrency 20–50.
Figure 22: Number of tools discovered per MCP server. MCP Bridge’s automatic tool discovery successfully enumerated all 50 tools across 4 servers upon initialization.
Figure 23: Latency percentile heatmap across all tool-execution operations. Cells report latency in milliseconds; darker shading indicates higher values.
Figure 24: Throughput–latency trade-off by concurrency level. Individual data points from 3 runs are shown alongside the connected mean trajectory, illustrating the trade-off between higher throughput and increased per-request latency.
Figure 25: Cold-start time breakdown. Total cold start for 4 MCP servers is 8,722 ms, dominated by Node.js module loading and server process initialization.
Figure 26: Multi-run reproducibility: mean latency per operation across 3 independent runs. Cross-run variance is low (std << 0.35 ms for all operations), confirming measurement stability.
Figure 27: Approximate cumulative distribution function (CDF) of latency per operation. Over 95% of requests complete within 3.5 ms for all operations.
Figure 28: API endpoint latency (mean and P95) for management endpoints (/health, /servers) and tool-listing endpoints. Management endpoints complete in under 1 ms.